Status Report

NASA Space Station On-Orbit Status 30 Jun 2004

By SpaceRef Editor
June 30, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 30 Jun 2004
http://images.spaceref.com/news/iss.11.jpg

All ISS systems continue to function nominally except those noted previously or below.   EVA-9B Day!!

CDR Gennady Padalka and FE/SO Mike Fincke woke up at 8:30am EDT to another strenuous day, which has sleep time scheduled tomorrow morning at 5:00am. 

Before breakfast, the crew completed the mandatory Russian pre- & post-EVA biomedical assessment MO-9/Urinalysis, and after breakfast, Padalka and Fincke tackled the previously exercised detailed list of pre-EVA closedown activities in the U.S. (USOS) and Russian (RS) segments to prepare the ISS for the six-hour unmanned period:

In the USOS, Mike reconfigured the ITCS (internal thermal control system) in the Lab for unmanned operation.   [[The procedure left string 1 of critical USOS flight electronics be cooled by the LTL (low temperature loop) pump and the string 2 cooled by the MTL (moderate temperature loop) pump.  This is done so that an MTL failure does not cause a complete loss of cooling to all USOS avionics.  To prevent or at least minimize any minor leakage past some of the QDs (quick disconnects), as seen in this configuration during the last EVA, Fincke today cycled (disconnected/reconnected) each QD once prior to demating it (to try flushing any precipitates that may have built up on the seals), and after disconnecting them he installed caps and plugs for the QDs that remain demated (to provide a second seal to leakage).]]

In the RS, Gennady set up the DC1 docking compartment and the SM PkhO (Service Module Transfer Compartment) for isolation and EVA activities.   [Preparations in the DC1 were mostly to confirm that all necessary gear is in place (e.g., fire extinguisher, pressure gauge, lighting fixtures, medical kits, etc.).]

Hatch closure between USOS and RS was at ~10:15am.   [Four hatches were closed, viz., Lab/Node, Node/Lab, Node stbd/Airlock, PMA-1/Node.]

Now locked out of the USOS, the crew has begun reconfiguring the RS for unmanned ops.   [In the SM, preparations include powering off the manual controls console (InPU) with its laptop, the pressure alarm sensors (DSD), the C&W (caution & warning) panel (PSS), the digital clock, and all laptops in the Service Module (SM). Gennady also closes the protective window covers and deactivates the TVIS treadmill, the running payloads, the food supply subsystem (SOP) with its food warmers, the condensate water processing system (SRV-K2), the thermal control system (SOTR) and #1 air conditioner (SKV-1), the toilet system (ASU), the Elektron O2 generator, the Vozdukh CO2 scrubber, and the ventilation system.]    

Later, the crewmembers perform a checkout of the Orlan systems and their BSS interface units in the DC1 and PkhO.  Testing of communication links follows, including VHF/voice and biomedical electrode belt and telemetry hookups via the BSS (later by the wireless in-suit Tranzit-B radio telemetry system) for vital signs and equipment monitoring. 

At ~2:00pm, Fincke removed the ventilation air duct extending from the SM through the DC-1 to the Soyuz TMA-4, including its interior fans, while Padalka checked out the four BK-3 oxygen repressurization tanks in the DC1.

Inside the DC-1, Mike and Gennady will then begin suit donning and backpack sealing at ~3:25pm. 

At 3:00pm, ISS attitude control will be handed over to RS MCS (motion control system), to be returned to the U.S. CMGs and momentum management mode at 10:50pm.  Other ground-commanded activities include: (1) MCC-H activating both PCUs (plasma contactor units) at 4:00pm, switching the BGAs (beta gimbal assemblies) of the U.S. solar array wings to Solar Autotrack mode, turning on the outside CETA (crew external translation aid) lights and configuring the MSS (mobile service system) external cameras, while (2) TsUP/Moscow activates “Tranzit-B” suit communications, deactivates the GTS (Global Timing System) transmitter and feathers the FGB solar arrays.  During the EVA, MCC-H will also conduct an on-orbit checkout plus 60-sec spin-up of CMG-2 (at 9:55pm) and switch the DDCU (dc-to-dc converter unit) 2A, i.e., the RPCM’s power source, first off, later back on.

After final checkout of suits and their controls, successive stages of airlock depressurization will begin at ~3:55pm, each step allowing for a test for Orlan leaks.   [Pressure inside the spacesuits will be reduced to 0.42 at (6.2 psi).  After suit purge, a 30-minute oxygen prebreathe period starts at ~5:00pm, as pressures between DC1 and the SM transfer compartment (PkhO) are equalized and then further reduced.  A final leak check will be conducted of the BK-3 O2 tanks.  At end of prebreathe, DC1 pressure will be down to 15 mmHg (torr), holding for 5 min for a final cabin leak check, followed by switch to independent Orlan suit power.]

EV1 hatch opening is expected at or about 5:40pm. 

After egress at ~5:46pm, the crew will configure and prepare the GStM1 “Strela” cargo crane at the DC1, extend it to its full length and attach its “business end” at an FGB handrail.  They will then translate along the boom to the PMA-1 (pressurized mating adapter 1) between FGB and Node, move to the S0 truss element on top of the Lab and perform the R&R (removal & replacement) of the defunct S0-2B RPCM (remote power controller module) with the spare.  After cleaning up the S0 worksite, the crew will return to the DC1, while reversing all previous activity steps, to ingress the airlock at or about 11:40pm tonight.

After repressurizing the DC1 airlock, the crew will open hatches and reenter the SM.  The RS cabin will be repressed with fresh air from Progress 14P, as per joint agreement, to a total pressure of 730-735 mmHg (Torr), followed later by a nitrogen (N2) repress of 19-20 mmHg from U.S. N2 tanks.

Immediately afterwards, Padalka and Fincke will take another MO-9 urine biochemistry test (~00:45am), followed by restoring comm systems to pre-EVA configuration.

Hatch opening to the U.S. segment is scheduled for ~1:45am, after which Mike Fincke will reconfigure the ITCS and the OpsLAN onboard computer network.  Meanwhile, Gennady Padalka will restore the DC1 and other RS modules to pre-EVA conditions.

Sleep time on the current offnominal shifted cycle begins around 5:00am tomorrow morning, to extend through 2:00pm tomorrow afternoon.

During the night of 6/29, the crew noticed unusual vibrations caused by an attitude disturbance due to off-nominal RS thruster firings during a test on a new Progress 14P angular rate sensor (DUS).  No impact on the EVA.   [During the test maneuver, which was kicked off by two Progress thrusters, one command confirmation message from the thrusters sent by the US-21 matching unit was lost, as has happened before on US-21s.  This was interpreted by the motion control system (MCS) as a thruster failure in manifold 2, whereupon it selected manifold 1 in addition to manifold 2.  With that, two additional Progress thrusters (from the second manifold) began to fire for pitch and yaw control.  Some time after this, the second confirmation message from the thrusters was also lost, causing the MCS to deactivate all Progress thrusters and select SM thrusters (two SM jets for yaw and pitch).  The combination of going from two to four thrusters and then back to two probably resulted in the stack vibrations reported by Mike.  Analysis continues.]

A temporary boot-up failure by an IBM ThinkPad A31p (NGSD-8) due to inoperative fan yesterday morning was cleared up by Fincke by cleaning the fan area and rebooting.   [As reported by Expedition 7 during post-flight debrief, the forward end of the Lab, where NGSD-8 is located, collects more dust than other locations.]

Today’s CEO (Crew Earth Observations) photo targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, except for the shutter closure and condensation-prevention plan (limited to 90 min. in 24 hours), were Caracas, Venezuela (this city has proved to be a challenging target for ISS crews so far.  Today’s pass was near-nadir pass in good light and fair weather.  As ISS approached the Venezuelan coast from the NW, the crew was to look for the city just inland from the coast in a small E-W valley), Las Vegas, Nevada (monsoon moisture is now making its way into the desert southwest of the US and is producing heavy afternoon clouds. However, this near-nadir pass, in late morning light, should have provided a good view of the Strip, with Lake Meade and Hoover Dam just a few seconds later), and Typhoon Tingting (DYNAMIC EVENT: This tropical cyclone has a large asymmetrical cloud pattern with good banding features and an intermittently visible eye structure in recent weather satellite imagery.  It remains a Category 2 system but continues to slowly strengthen as it moves northward over Iwo Jima.  With a near-nadir pass, researchers hope for long lens views into the eye of the storm, if present. Otherwise, they would have liked oblique views of the banding structures shot with the 180mm lens or perhaps obliques on approach and departure).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. & Russian Segment Status  (as of today, 1:49pm EDT)

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is On.  Vozdukh CO2 scrubber is On.  U.S. CDRA CO2 scrubber is Off.  TCCS (trace contaminant control subsystem) is operating.  SM Gas Analyzer has been calibrated and is used for ppO2 & ppCO2 monitoring.  MCA (major constituents analyzer) is in Life Extending Mode (LEM).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On, SKV-2 is Off (SM panel mods completed 4/8; SKV-2 activation failed 4/20; is now functioning again).  SFOG slot #2 fan suspect (not usable).
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 23.6; ppO2 (mmHg) — 150.7; ppCO2 (mmHg) — 5.3.
  • SM Transfer Compartment:  Pressure (mmHg) — 743; temperature (deg C) — 20.3.
  • FGB Cabin:  Pressure (mmHg) — 755; temperature (deg C) — 20.8.
  • Node:  Pressure (mmHg) — 751.5; temperature (deg C) — 23.1 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 753.5; temperature (deg C) — 25.0; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 753.3; temperature (deg C) — 23.4; shell heater temp (deg C) — n/a, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • (n/a = data not available)
  • PMA-1:  Shell heater temp (deg C) — n/a
  • PMA-2:  Shell heater temp (deg C) — n/a.

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in directed position (Dual angle/”blind” mode, non solar-tracking, biased for drag reduction).
  • SM batteries:  All batteries (8) are on line in “Partial Charge” mode.  
  • FGB batteries:  Battery #5 is off line; all other batteries (5) are on line in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Standby mode (will both be in Discharge for the EVA).

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is backup, and C&C-3 is in standby.
  • GNC-2 MDM is prime; GNC-1 is backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-1 is On (primary), EXT-2 is Off (backup).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22/03).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22/03).
  • FGB MDM-1 is Off (failed, 11/21/03); MDM-2 is Operational.

Propulsion System:

  • Total propellant load available: 3963 kg (8737 lb) as of 6/24/04;  [SM(552) + FGB(2772) + Progress M(639)].  (Capabilities: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 2 CMGs on-line (CMG-1 failed, since 6/6/02; CMG-2’s RPC-17 failed 4/21/04).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH XVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch: -9 deg, roll: 0 deg]), with CMG/TA (thruster assist) Momentum Management, until 6/28, following the EVA.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally (may require a mask).
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.

Robotics:

  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, operational on redundant string, off on prime.
  • MBS: KA (keep alive) power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 8:18am EDT [= epoch]):

  • Mean altitude — 360.5 km
  • Apogee — 364.2 km
  • Perigee — 356.8 km
  • Period — 91.7 min.
  • Inclination (to Equator) —  51.6319 deg
  • Eccentricity — 0.0005451
  • Solar Beta Angle — 25.0 deg
  • Orbits per 24-hr. day — 15.69
  • Mean altitude loss in last 24 hours — 80 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 32025

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see

 http://www.hq.nasa.gov/osf/station/viewing/issvis.html

SpaceRef staff editor.